OpenStudy (anonymous):

Find all values of x for which the function is differentiable for y = sin x - x cos x.

OpenStudy (agent47):

y'=Cosx-(x*(-Sinx)+Cosx)

OpenStudy (agent47):

y' = Cosx + x*Sinx - Cosx = x*Sinx

OpenStudy (agent47):

that's a continuous function, so y is differentiable for all x

OpenStudy (anonymous):

Okay, thank you so much!

OpenStudy (anonymous):

How did you find y' though?

OpenStudy (agent47):

y' = derivative of y

OpenStudy (agent47):

y = sin x - x cos x. Derivative of Sin x is Cos x To find the derivative of x*Cosx, use the product rule

OpenStudy (anonymous):

Oh okay. Thanks!

OpenStudy (agent47):

(f*g)'=f*g'+g*f'

OpenStudy (agent47):

and that's the product rule if ur wondering

OpenStudy (anonymous):

Thanks :).

OpenStudy (anonymous):

How would you do it for y = \[\sqrt{(x-1)/(1+x ^{2})}\]